Which of the following is an equivalent expression to sin θ − (sin θ)(cos2θ)?


1
Expert's answer
2022-05-01T16:06:58-0400
"\sin \theta-(\sin\theta)(\cos^2\theta)=\sin \theta(1-\cos^2\theta)"

"=\sin \theta(\sin^2 \theta)=\sin^3 \theta"

"\sin \theta-(\sin\theta)(\cos^2\theta)=\sin^3 \theta"